loughamy waketech edu My Turbo Tax Login Signint x f Me 2019 Connect Required: Ch 2 Quiz Saved Help Save & Exit Submit Dehner Corporation uses a job-order costing system with a single plantwide predetermined overhead rate based on direct labor-hours. The company based its predetermined overhead rate for the current year on the following data: points Total direct labor-hours Total fixed manufacturing overhead cost Variable manufacturing overhead per direct labor-hour 02-) Recently, Job P951 was completed with the following characteristics: Number of units in the job Total direct labor-hours Direct materials Direct labor cost 25 100 $840 $10, 100 The total job cost for Job P951 is closest to: (Round your intermediate calculations to 2 decimal places.) Multiple Choice $1.730 $10,990 Mc Graw Hait くPree 20f 9llǐ Next>

Connect Required: Ch 2 Quiz Saved Help Save & Exit Submit Variable manufacturing overhead per direct labor-hour 2 Recently, Job P951 was completed with the following characteristics: points Number of units in the 25 Total direct labor-hours Direct materials Direct labor cost 10θ $ 840 $10,100 The total job cost for Job P951 is closest to: (Round your intermediate calculations to 2 decimal places.) Multiple Choice $1,730 $10,990 $10,940 $11,830 Mc Graw HiH https

Total direct labor-hours:101,000 Total fixed manufacturing overhead cost: $595,900 Variable manufacturing overhead per direct labor-hour: $3.00
Answer #1

$ 11830 is the Correct Answer

Calculations :

CALCULATE TOTAL MANUFACTURING COST FOR JOB P951:

PARTICULARS AMOUNT
DIRECT MATERIAL (Given ) $840
DIRECT LABOR COST (Given) $10100
VARIABLE MANUFACTURING OVERHEAD (100 DLH * $3 PER DLH) (Note 1) $300
FIXED MANUFACTURING OVERHEAD (595900*100/101000) (Note 2) $590
TOTAL MANUFACTURING COST For Job P951 ( Total of above items) $11830

(Note 1) VARIABLE MANUFACTURING OVERHEAD= Total hours used in Job P951 * Variable manufacturing overhead per direct Labor-hour =  = 100 * 3 = $300   

(Note 2) FIXED MANUFACTURING OVERHEAD=(fixed manufacturing overhead cost * Direct Lobor hours used in Job P951)/Total  Direct Labor-Hours = = 595900*100/101000 =  $590

DLH Means = Direct Labor-Hours